OpenStudy (anonymous):

let v1=(2,-6) and v2= (-4,7) (a) compute |v1| and |v2|

OpenStudy (nurali):

x1=2 y1=-6 \[ |v1|=\sqrt{x1^2+y1^2}\]

OpenStudy (nurali):

can u solved v1 now?

OpenStudy (anonymous):

yes i got sqrt -36

OpenStudy (anonymous):

*32

OpenStudy (nurali):

wrong

OpenStudy (anonymous):

what did I do wrong then?

OpenStudy (anonymous):

sqrt of 40

OpenStudy (nurali):

\[|v1|=\sqrt{(2)^2+(-6)^2}\]

OpenStudy (nurali):

correct

OpenStudy (nurali):

can u solved v2 now?

OpenStudy (anonymous):

i got sqrt of 40 I messed up with the negative sign

OpenStudy (anonymous):

how would i find |v2| ?

OpenStudy (nurali):

x1=-4 y1=7

OpenStudy (nurali):

can u solved v2 now?

OpenStudy (anonymous):

sqrt of 65. How would I also do v1+v2? would i justadd 2-4 and multiply it by -6+7?

OpenStudy (nurali):

v1+v2=2x1-6y1-4x1+7y1=-2x1+y1=(-2,1)